# Model-Based Lookahead Reinforcement Learning for in-hand manipulation

### v1

@@ I Introduction
While very complex robotic hands that could be used as a substitute for the grippers already exist, such as the Shadow Hand [^1] https://www.shadowrobot.com/products/dexterous-hand/, using them comes with another challenge: the system needs to learn how to use the hand to manipulate objects. A way to do this is to use a method that takes inspiration from how humans learn, Reinforcement Learning (RL). Although RL explain the main mechanism behind reward learning, execution of real-time tasks by humans requires additional components such as sensorial anticipation, in order to deal with latencies [1]. These anticipatory mechanisms rely on internal simulations of the forthcoming sensorial input, which works like the Model Predictive Control (MPC) approach [2, 3]. From the RL point of view, MPC corresponds use Model-Based RL (MBRL) in a short-time horizon. Since Model-Based requires an accurate model of the dynamics of the environment, it is very difficult to apply it to in-hand object manipulation. Thus, we follow the hybrid model that combines Model-Free RL (MFRL) and MBRL. One of these methods is Model-Based Lookahead Reinforcement Learning [4]. The main focus of this work is to apply this method with robotics, and conclude if the promising results achieved in the previous tests are still achievable for in-hand manipulations tasks using a dexterous robotic hand.
@@ II-A Introduction of Notation
RL problems can be described as Markov Decision Processes [5], using the tuple $\{S,A,R,T,\gamma\}$. Here, $S$ represents the state space, and $A$ the action space. $R(s,a)$ represents a reward function, which returns the immediate reward for taking an action $a$ in a state $s$. $T(s^{\prime}|s,a)$ is the dynamics of the system, which returns the probability distribution of the next states $s^{\prime}$ after taking action $a$ in state $s$. Finally $\gamma\in[0,1]$ is a discount factor, applied to the sum of immediate rewards $G$, where $G=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})$. The objective of RL is to obtain a policy, $\pi$ (i.e. controller in our case) that selects actions in order to maximize the rewards.
@@ II-B Model-Free Reinforcement Learning
In cases such as in-hand manipulation of objects, traditional models of the dynamics require to know all contact points and forces. These forces can be computed only in the case of rigid objects, but robotic hands are usually built with elastic components. Thus, the dynamics model $T$ suffer from a large reality gap that limits the application of purely MPC models. MFRL approaches the problem by learning policies from exploration, without having to learn a model of its dynamics. Two ways of approaching MFRL are policy-based and value-based methods: (i) In value function methods, the objective is to learn the value of the states $V$, or the value of state-action pairs $Q$, that is, the expected rewards obtained when following a policy $\pi$, starting at a certain state/state-action pair. The optimal value function $V^{*}(s)$ is defined as $V^{*}(s)=\max_{\pi}\mathbb{E}\left[G_{t}\mid s_{t}=s\right]$, and the optimal action-value function $Q^{*}(s,a)$ is defined as $Q^{*}(s,a)=\max_{\pi}\mathbb{E}\left[G_{t}\mid s_{t}=s,a_{t}=a,\pi\right]$. (ii) In policy-based methods the objective is to parameterize a policy. The policy $\pi$ will depend on a parameter vector $\theta$. This approach is based on the policy gradient theorem, which gives a proportional equation to the gradient of the performance measure, $J(\theta)$, updating the parameter vector as $\theta_{t+1}=\theta_{t}+\alpha\nabla J(\boldsymbol{\theta})$. The gradient of the performance measure is derived in [5].
We explore the use of an Actor-Critic method, which uses both a value function (critic) and a policy parametrization (actor). While the actor decides on which action to take, the critic will evaluate the action-state pair and give the actor a measurement of how good the action is. These algorithms, such as PPO [6] and SAC [7] have had success in in-hand manipulation tasks, such as [8], where the objective was to manipulate a cube with labeled sides to the correct configuration using vision.
@@ II-C Model-Based Reinforcement Learning
A different approach to MFRL which is more data efficient [9] is MBRL. By interacting with the environment, MBRL learns a model which represents the dynamics of the system. After learning, the model is then used for planning. Planning can be done offline, where the algorithm uses the information from the model to obtain a policy, or online, such as with MPC. MPC has shown very good results at selecting the right action while considering a small (i.e. non-infinite) horizon. Its sucess comes from very good model of the dynamics of the environment. While MBRL learning models require less data then MFRL, they require a very accurate dynamics model. This dynamics model can be modeled as deterministic or probabilistic. Deterministic models assume that the environment is predictable and without randomness or noise, meaning that for the same state and action the output will always be the same, $s^{\prime}=f(s,a)$, where as probabilistic models output the probability distribution of the system dynamics $T(s^{\prime}|s,a)$. In robotic tasks, success has been found in using deep NNs for the dynamics model. In [10], a robot using a knife is used to cut different foods. The NN model is then applied online with the use of an MPC controller.
@@ II-D Combining MFRL and MBRL
While current efforts study how to develop a generalist world model (i.e. a multiple task that represents the dynamics as latent representation) that combines MFRL and MBRL, these methods require very large computational resources [11, 12]. We focus on previous works that (i) are task specific and develop models with (ii) relatively small number of parameters [13, 14, 15, 4]. These works leverage the strengths of MFRL and MBRL in a hybrid manner, such as the Model-based Lookahead Reinforcement Learning [4]. We select [4] because the action selection approach is not as greedy as in [13], and has less parameters to tune than in [14, 15]. In [4], the authors propose that, after training a policy and value-function through an Actor-Critic Method, and a dynamic model with the extracted datasets from training, trajectory evaluation is then applied during online planning, as done with an MPC. This approach will serve as the base method for this work. The model is divided in two parts: the training phase, and the evaluation phase: The training phase begins by training in parallel with the same data, the policy $\pi_{\theta^{\pi}}$, the value function $V_{\theta^{V}}$ and the transition model $f_{\theta^{f}}$. All of these are trained as NNs, as parametric functions with parameters $\theta^{\pi}$$\theta^{V}$,and $\theta^{f}$ respectively.
@@
After evaluation, the trajectory with the highest reward could be chosen, and its first action would be used, but this might be too optimistic [4], as the simulated rewards are obtained using a learned dynamics model and policy, and not an actual model of the environment. A way to reduce the approximation errors from this method is to, instead of greedily choosing the best trajectory, using the average of the best $E$ trajectories, where $E$ is a fixed chosen value. This is performed by using a sorting algorithm, and finally outputting the averaged first action of the best trajectories.
@@
Unlike most approaches in which there is only one NN to train, this method needs to train three separate NNs. Even if the model has a high sample efficiency reducing the size of the needed data set, the computational resources (such as number of GPU tensor cores and available memory) needed to train the networks in parallel is much higher, especially in highly complex problems such as in-hand manipulation. Another issue is the high amount of hyperparameters. Fine tuning these values will be time consuming, and will require more experimentation. Finally, as mentioned in [4], the results obtained from using horizon values 5 and 20 are very similar. As the trajectories are obtained from an approximated dynamics model, they will accumulate errors, a problem that may result in poor prediction of trajectories for very large horizons. This may hinder performance, as longer horizons allow for more extensive exploration [13].
@@ III Methodology
On in-hand manipulation, many tasks require access to the position or orientation of an object (Ex: rotating a cube to a labeled face [8]). With the current information the dexterous hand provides, this is not possible, as it only provides access to the feature values of the hand and its contact points. A way to obtain the pose of the objects, is to use a pose detector, which can be substituted in most physic simulators. Because of this, it is a design choice to fully develop this work in a simulated environment.
@@ III-A Dynamics
The simulator used in this work is Nvidia Isaac Gym [16]. This simulator provides a physics engine, as well as integration with various frameworks such as SKRL [17], a python library for RL which is integrated with Nvidia Isaac Gym, the PyTorch framework [18], and Isaac Gym Benchmark Environments (IsaacGymEnvs) [19], a repository containing various tasks, objects and other assets. Another important feature of Isaac Gym is the use of multi-environments, which will increase the number of training samples. The models will be trained with $2^{13}$ parallel environments. The simulator gives access to the state vectors of the simulated objects over each timestep, such as position, orientation, velocity and acceleration, where the position, linear velocity and linear acceleration are given as a three dimensional vector $(x,y,z)$ and the orientation is defined as quaternions $(q_{1},q_{2},q_{3},q_{4})$, removing the need for a pose detector.
@@
To perform this in Nvidia Isaac Gym, this work uses a system based on the work [20] which, given the position and linear velocity of the finger joints, as well as the action variables, can calculate based on a spring coefficient matrix $K$ and Synergy Coefficient matrix $S$ the torque value of those finger joints. Since all of the DoF are revolute, the inputs are torque values in $Nm$. The rest of the joints are all controlled in position, which are the wrist joints and thumb-palm joint, in radians.
@@ III-B Manipulation Task
We focus on consecutively rotating an object on the palm of the hand to goal orientations, for a fixed period of time. This task is obtained from the Isaac Gym Benchmark Environments, which is based on the task [8] from OpenAI. The task is obtained from utilizing two instances of the same object: object, and goal object. For a defined maximum episode time, the object is spawned on top of the palm of the hand on a random orientation, where the goal is to be inside a small range off the orientation of the goal object, which is spawned in the simulation at the same time, with a random orientation as well. If the object matches the orientation within the defined maximum episode time (600 timesteps), the goal object is now re-spawned with a new orientation, while the object continues in the same position, and must match the new orientation. The reset condition will be if the object has fallen from the palm, represented by a defined fall distance parameter. The number of consecutive times the object matches the orientation of the goal object without failing will be recorded, as well as the average episode time. The task will be performed for three different objects: a cube, an egg, and a parallelepiped. The three objects have varying geometrical complexities which will result in tasks with different difficulties, which may give more insight on the performance gain of Model-Based Lookahead RL.
@@ IV-A Obtaining the Policy
An interesting result is the better performance of the egg compared to the cube, as the egg has a much more complex geometry which leads to more complex interactions between the hand and object. A possibility for this result may come from the fact that the policy, in contrast to results such as [8], during manipulation the robotic hand tends to roll the object on the palm, assisted by the finger to rotate them, instead of picking up the object with opposing fingers, holding it with the thumb. In the obtained results of this work, the thumb is barely used, only applying forces on the object without much result in movement. By lowering the pitch DoF of the palm, the robot rotates the object with the pinky, ring, middle and index fingers on the palm instead of picking it up, which leads to better results when the objects have rounder surfaces, such as the egg, as opposed to the cube and parallelepiped. Since the cube has a smaller and uniform size, even with sharp edges, the robot still manages to rotate it with ease, however, the addition of different sizes for the x axis of the parallelepiped (the x axis is double the value of y and z, which are the same size as the cube) results in harder movements for the robot to perform, reducing the overall performance significantly.
@@ IV-B Obtained the Dynamics Model
Having the extracted the datasets from each task, it is now possible to train the dynamic models of each one. The dynamic models with highest performance metrics obtained utilize the finger joints and objects position and linear velocity. Since position control utilizes both position and velocity errors in the PD controller, these features are essential for obtaining low prediction errors for the finger joints. For training the neural network, Supervised Learning is performed, aided with the Adam optimizer [21]. The configuration used repeats a block containing a Linear Fully Connected Layer, followed by Batch Normalization, a ReLU activation function and a dropout layer with probability of 20%. This is enough for the under-actuated system, however the for the fully-actuated system due to the large action-space, a modular approach is added, using an additional network (Single Joint Model) at the start which takes as input the joint position, velocity, and action, and outputs the predicted position and velocity. This is done for each joint, as each joint is independent from each other, simplifying the prediction of the joint values, which are then used as input in the Dynamic Model. The results of these network are shown in tables III and IV:
@@ IV-C Model-based Lookahead Reinforcement Learning
Having obtained the policy and value-function from training a PPO algorithm, and training a dynamic model from the datasets extracted for all objects with high reward/minimal error, it is now possible to combine them in evaluation mode as explained in section II-D. As trajectory evaluation is referred as MFRL-MPC [4], for further tests the model using PPO as base for the policy will be referred as PPO-MPC.
@@ IV-F Under-Actuated System
Unlike the results obtained in the fully-actuated system, in the under-actuated system the improvement gained from trajectory evaluation is almost negligible in the cube and parallelepiped tasks, but noticeable in the egg task, with an increase of 0.6 Consecutive successes. Overall, besides the egg task, the results of both cube and parallelepied fail to obtain any significant improvement, even though the obtained dynamic models have similar losses to the dynamic models obtained in the fully-actuated system. Due to the restrictions of the joints, the policies obtained have much smaller Average Rewards and number of Consecutive Successes, as turning the objects to a precise orientation becomes much more difficult with limited movements. Similar to the results of V, with the Reward Function that uses differences in orientation and action penalty smoothing used for training, policies with higher Average Reward values seem to obtain the largest improvement from trajectory evaluation, which is also similar to overall results obtain in [4]. As such, when it is not possible to obtain a good policy for a system, the performance gain from Model-Based Lookahead RL is minimal, even if the obtained dynamic model has high accuracy. Since this under-actuated simulation aims to mimic the real-world RH8DR hand, it is likely that in the real system the benefits of trajectory evaluation are also minimal.

### v2

@@ I Introduction
While very complex robotic hands that could be used as a substitute for the grippers already exist, such as the Shadow Hand [^1] https://www.shadowrobot.com/products/dexterous-hand/, using them comes with another challenge: the system needs to learn how to use the hand to manipulate objects. A way to do this is to use a method that takes inspiration from how humans learn, Reinforcement Learning (RL). Although RL explain the main mechanism behind reward learning, execution of real-time tasks by humans requires additional components such as sensorial anticipation, in order to deal with latencies [WOLPERT1998338]. These anticipatory mechanisms rely on internal simulations of the forthcoming sensorial input, which works like the Model Predictive Control (MPC) approach [mpc-1982, RICHALET19931251]. From the RL point of view, MPC corresponds use Model-Based RL (MBRL) in a short-time horizon. Since Model-Based requires an accurate model of the dynamics of the environment, it is very difficult to apply it to in-hand object manipulation. Thus, we follow the hybrid model that combines Model-Free RL (MFRL) and MBRL. One of these methods is Model-Based Lookahead Reinforcement Learning [hong2019model]. The main focus of this work is to apply this method with robotics, and conclude if the promising results achieved in the previous tests are still achievable for in-hand manipulations tasks using a dexterous robotic hand.
@@ II-A Introduction of Notation
RL problems can be described as Markov Decision Processes [sutton2018reinforcement], using the tuple $\{S,A,R,T,\gamma\}$. Here, $S$ represents the state space, and $A$ the action space. $R(s,a)$ represents a reward function, which returns the immediate reward for taking an action $a$ in a state $s$. $T(s^{\prime}|s,a)$ is the dynamics of the system, which returns the probability distribution of the next states $s^{\prime}$ after taking action $a$ in state $s$. Finally $\gamma\in[0,1]$ is a discount factor, applied to the sum of immediate rewards $G$, where $G=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})$. The objective of RL is to obtain a policy, $\pi$ (i.e. controller in our case) that selects actions in order to maximize the rewards.
@@ II-B Model-Free Reinforcement Learning
In cases such as in-hand manipulation of objects, traditional models of the dynamics require to know all contact points and forces. These forces can be computed only in the case of rigid objects, but robotic hands are usually built with elastic components. Thus, the dynamics model $T$ suffer from a large reality gap that limits the application of purely MPC models. MFRL approaches the problem by learning policies from exploration, without having to learn a model of its dynamics. Two ways of approaching MFRL are policy-based and value-based methods: (i) In value function methods, the objective is to learn the value of the states $V$, or the value of state-action pairs $Q$, that is, the expected rewards obtained when following a policy $\pi$, starting at a certain state/state-action pair. The optimal value function $V^{*}(s)$ is defined as $V^{*}(s)=\max_{\pi}\mathbb{E}\left[G_{t}\mid s_{t}=s\right]$, and the optimal action-value function $Q^{*}(s,a)$ is defined as $Q^{*}(s,a)=\max_{\pi}\mathbb{E}\left[G_{t}\mid s_{t}=s,a_{t}=a,\pi\right]$. (ii) In policy-based methods the objective is to parameterize a policy. The policy $\pi$ will depend on a parameter vector $\theta$. This approach is based on the policy gradient theorem, which gives a proportional equation to the gradient of the performance measure, $J(\theta)$, updating the parameter vector as $\theta_{t+1}=\theta_{t}+\alpha\nabla J(\boldsymbol{\theta})$. The gradient of the performance measure is derived in [sutton2018reinforcement].
We explore the use of an Actor-Critic method, which uses both a value function (critic) and a policy parametrization (actor). While the actor decides on which action to take, the critic will evaluate the action-state pair and give the actor a measurement of how good the action is. These algorithms, such as PPO [schulman2017proximal] and SAC [haarnoja2018soft] have had success in in-hand manipulation tasks, such as [andrychowicz2020learning], where the objective was to manipulate a cube with labeled sides to the correct configuration using vision.
@@ II-C Model-Based Reinforcement Learning
A different approach to MFRL which is more data efficient [deisenroth2013survey] is MBRL. By interacting with the environment, MBRL learns a model which represents the dynamics of the system. After learning, the model is then used for planning. Planning can be done offline, where the algorithm uses the information from the model to obtain a policy, or online, such as with MPC. MPC has shown very good results at selecting the right action while considering a small (i.e. non-infinite) horizon. Its sucess comes from very good model of the dynamics of the environment. While MBRL learning models require less data then MFRL, they require a very accurate dynamics model. This dynamics model can be modeled as deterministic or probabilistic. Deterministic models assume that the environment is predictable and without randomness or noise, meaning that for the same state and action the output will always be the same, $s^{\prime}=f(s,a)$, where as probabilistic models output the probability distribution of the system dynamics $T(s^{\prime}|s,a)$. In robotic tasks, success has been found in using deep NNs for the dynamics model. In [lenz2015deepmpc], a robot using a knife is used to cut different foods. The NN model is then applied online with the use of an MPC controller.
@@ II-D Combining MFRL and MBRL
While current efforts study how to develop a generalist world model (i.e. a multiple task that represents the dynamics as latent representation) that combines MFRL and MBRL, these methods require very large computational resources [Hansen2022tdmpc, hansen2024tdmpc2]. We focus on previous works that (i) are task specific and develop models with (ii) relatively small number of parameters [lowrey2018plan, bhardwaj2021blending, bejjani2021learning, hong2019model]. These works leverage the strengths of MFRL and MBRL in a hybrid manner, such as the Model-based Lookahead Reinforcement Learning [hong2019model]. We select [hong2019model] because the action selection approach is not as greedy as in [lowrey2018plan], and has less parameters to tune than in [bhardwaj2021blending, bejjani2021learning]. In [hong2019model], the authors propose that, after training a policy and value-function through an Actor-Critic Method, and a dynamic model with the extracted datasets from training, trajectory evaluation is then applied during online planning, as done with an MPC. This approach will serve as the base method for this work. The model is divided in two parts: the training phase, and the evaluation phase: The training phase begins by training in parallel with the same data, the policy $\pi_{\theta^{\pi}}$, the value function $V_{\theta^{V}}$ and the transition model $f_{\theta^{f}}$. All of these are trained as NNs, as parametric functions with parameters $\theta^{\pi}$$\theta^{V}$,and $\theta^{f}$ respectively.
@@
After evaluation, the trajectory with the highest reward could be chosen, and its first action would be used, but this might be too optimistic [hong2019model], as the simulated rewards are obtained using a learned dynamics model and policy, and not an actual model of the environment. A way to reduce the approximation errors from this method is to, instead of greedily choosing the best trajectory, using the average of the best $E$ trajectories, where $E$ is a fixed chosen value. This is performed by using a sorting algorithm, and finally outputting the averaged first action of the best trajectories.
@@
Unlike most approaches in which there is only one NN to train, this method needs to train three separate NNs. Even if the model has a high sample efficiency reducing the size of the needed data set, the computational resources (such as number of GPU tensor cores and available memory) needed to train the networks in parallel is much higher, especially in highly complex problems such as in-hand manipulation. Another issue is the high amount of hyperparameters. Fine tuning these values will be time consuming, and will require more experimentation. Finally, as mentioned in [hong2019model], the results obtained from using horizon values 5 and 20 are very similar. As the trajectories are obtained from an approximated dynamics model, they will accumulate errors, a problem that may result in poor prediction of trajectories for very large horizons. This may hinder performance, as longer horizons allow for more extensive exploration [lowrey2018plan].
@@ III Methodology
On in-hand manipulation, many tasks require access to the position or orientation of an object (Ex: rotating a cube to a labeled face [andrychowicz2020learning]). With the current information the dexterous hand provides, this is not possible, as it only provides access to the feature values of the hand and its contact points. A way to obtain the pose of the objects, is to use a pose detector, which can be substituted in most physic simulators. Because of this, it is a design choice to fully develop this work in a simulated environment.
@@ III-A Dynamics
The simulator used in this work is Nvidia Isaac Gym [liang2018gpu]. This simulator provides a physics engine, as well as integration with various frameworks such as SKRL [serrano2023skrl], a python library for RL which is integrated with Nvidia Isaac Gym, the PyTorch framework [NEURIPS2019_9015], and Isaac Gym Benchmark Environments (IsaacGymEnvs) [makoviychuk2021isaac], a repository containing various tasks, objects and other assets. Another important feature of Isaac Gym is the use of multi-environments, which will increase the number of training samples. The models will be trained with $2^{13}$ parallel environments. The simulator gives access to the state vectors of the simulated objects over each timestep, such as position, orientation, velocity and acceleration, where the position, linear velocity and linear acceleration are given as a three dimensional vector $(x,y,z)$ and the orientation is defined as quaternions $(q_{1},q_{2},q_{3},q_{4})$, removing the need for a pose detector.
@@
To perform this in Nvidia Isaac Gym, this work uses a system based on the work [catalano2014adaptive] which, given the position and linear velocity of the finger joints, as well as the action variables, can calculate based on a spring coefficient matrix $K$ and Synergy Coefficient matrix $S$ the torque value of those finger joints. Since all of the DoF are revolute, the inputs are torque values in $Nm$. The rest of the joints are all controlled in position, which are the wrist joints and thumb-palm joint, in radians.
@@ III-B Manipulation Task
We focus on consecutively rotating an object on the palm of the hand to goal orientations, for a fixed period of time. This task is obtained from the Isaac Gym Benchmark Environments, which is based on the task [andrychowicz2020learning] from OpenAI. The task is obtained from utilizing two instances of the same object: object, and goal object. For a defined maximum episode time, the object is spawned on top of the palm of the hand on a random orientation, where the goal is to be inside a small range off the orientation of the goal object, which is spawned in the simulation at the same time, with a random orientation as well. If the object matches the orientation within the defined maximum episode time (600 timesteps), the goal object is now re-spawned with a new orientation, while the object continues in the same position, and must match the new orientation. The reset condition will be if the object has fallen from the palm, represented by a defined fall distance parameter. The number of consecutive times the object matches the orientation of the goal object without failing will be recorded, as well as the average episode time. The task will be performed for three different objects: a cube, an egg, and a parallelepiped. The three objects have varying geometrical complexities which will result in tasks with different difficulties, which may give more insight on the performance gain of Model-Based Lookahead RL.
@@ IV-A Obtaining the Policy
An interesting result is the better performance of the egg compared to the cube, as the egg has a much more complex geometry which leads to more complex interactions between the hand and object. A possibility for this result may come from the fact that the policy, in contrast to results such as [andrychowicz2020learning], during manipulation the robotic hand tends to roll the object on the palm, assisted by the finger to rotate them, instead of picking up the object with opposing fingers, holding it with the thumb. In the obtained results of this work, the thumb is barely used, only applying forces on the object without much result in movement. By lowering the pitch DoF of the palm, the robot rotates the object with the pinky, ring, middle and index fingers on the palm instead of picking it up, which leads to better results when the objects have rounder surfaces, such as the egg, as opposed to the cube and parallelepiped. Since the cube has a smaller and uniform size, even with sharp edges, the robot still manages to rotate it with ease, however, the addition of different sizes for the x axis of the parallelepiped (the x axis is double the value of y and z, which are the same size as the cube) results in harder movements for the robot to perform, reducing the overall performance significantly.
@@ IV-B Obtained the Dynamics Model
Having the extracted the datasets from each task, it is now possible to train the dynamic models of each one. The dynamic models with highest performance metrics obtained utilize the finger joints and objects position and linear velocity. Since position control utilizes both position and velocity errors in the PD controller, these features are essential for obtaining low prediction errors for the finger joints. For training the neural network, Supervised Learning is performed, aided with the Adam optimizer [kingma2014adam]. The configuration used repeats a block containing a Linear Fully Connected Layer, followed by Batch Normalization, a ReLU activation function and a dropout layer with probability of 20%. This is enough for the under-actuated system, however the for the fully-actuated system due to the large action-space, a modular approach is added, using an additional network (Single Joint Model) at the start which takes as input the joint position, velocity, and action, and outputs the predicted position and velocity. This is done for each joint, as each joint is independent from each other, simplifying the prediction of the joint values, which are then used as input in the Dynamic Model. The results of these network are shown in tables III and IV:
@@ IV-C Model-based Lookahead Reinforcement Learning
Having obtained the policy and value-function from training a PPO algorithm, and training a dynamic model from the datasets extracted for all objects with high reward/minimal error, it is now possible to combine them in evaluation mode as explained in section II-D. As trajectory evaluation is referred as MFRL-MPC [hong2019model], for further tests the model using PPO as base for the policy will be referred as PPO-MPC.
@@ IV-F Under-Actuated System
Unlike the results obtained in the fully-actuated system, in the under-actuated system the improvement gained from trajectory evaluation is almost negligible in the cube and parallelepiped tasks, but noticeable in the egg task, with an increase of 0.6 Consecutive successes. Overall, besides the egg task, the results of both cube and parallelepied fail to obtain any significant improvement, even though the obtained dynamic models have similar losses to the dynamic models obtained in the fully-actuated system. Due to the restrictions of the joints, the policies obtained have much smaller Average Rewards and number of Consecutive Successes, as turning the objects to a precise orientation becomes much more difficult with limited movements. Similar to the results of V, with the Reward Function that uses differences in orientation and action penalty smoothing used for training, policies with higher Average Reward values seem to obtain the largest improvement from trajectory evaluation, which is also similar to overall results obtain in [hong2019model]. As such, when it is not possible to obtain a good policy for a system, the performance gain from Model-Based Lookahead RL is minimal, even if the obtained dynamic model has high accuracy. Since this under-actuated simulation aims to mimic the real-world RH8DR hand, it is likely that in the real system the benefits of trajectory evaluation are also minimal.